# Segmentation-free Goodness of Pronunciation

### v2

@@ IV-C Pre-training and fine-tuning of the acoustic models
The “DNN” and “TDNN” acoustic models are trained from scratch using the kaldi recipes444./egs/librispeech/s5/local/{nnet2/run_5a_clean_100, nnet3/run_tdnn}.sh. All the other models are based on the wav2vec2.0-large-xlsr-53 [conneau21_interspeech] which is available on Huggingface [^5] https://huggingface.co/blog/fine-tune-xlsr-wav2vec2 and have a transformer-based structure [w2v2]. This model was fine-tuned on the Librispeech “train-100-clean” set according to the corresponding loss functions:
@@
  • enCTC: CTC with entropy regularization [CTC-ME]
  • esCTC: CTC with equal space constraint [CTC-ME]
Fine-tuning was performed for at most 10 epochs or by early stopping based on the validation loss.
@@ IV-D Evaluation
To evaluate the performance of MDD on CMU Kids, we opt for the threshold-free criterion area under the receiver operative characteristic curve (AUC-ROC). This is because we want a robust method to evaluate the performance of our GOP definitions that is not dependent on a specific threshold chosen to perform the MDD task. This also allows us to use all the data for testing because we do not need to set aside a training set for threshold optimization. Moreover, AUC-ROC is considered to be robust to label bias which is the case for our data.
On the other hand, for the MDD-oriented dataset Speechocean762 the task is to predict three different classes. In this case, we train and evaluate several typical MDD models using the standard training and test splits. We follow the recommendations of using Pearson Correlation Coefficient (PCC) from the dataset’s paper [speechocean762] as MDD performance metric.
To give a reference on the quality of the different acoustic models in this study, we also report on phoneme recognition results measured with phoneme error rate (PER). The recognition is based on a naïve beam-search decoder without language models.
@@ IV-E Experiments
The set of experiments we include in this paper aim at answering the following questions:
1) Which GOP definition results in the best MDD performance? In this case we test GOP-X-Avg, GOP-X-SA, and the different variants of GOP-X-AF (GOP-X-AF-SDI, GOP-X-AF-SD, GOP-X-AF-S) where “X” corresponds to acoustic models trained with different loss functions and architectures as explained in Sec.IV-C.
2) How are the results dependent on the peakiness of the ASR models used for the assessment? In this case we test the different GOP definitions in combination with models that are trained with loss functions that are specifically defined to control peakiness: CTC, EnCTC, EsCTC and CE. We use our best method GOP-AF-SD as alignment-free version in this experiment and due to numerical problem with CE-trained models, we use the “numerical” implementation as discussed in Sec.III-B.
3) How are the results for GOP-AF dependent on the length of the context around the target phoneme? In this case we create utterances by cropping the original recordings and gradually increasing the context around the target phoneme. Each time we add one phoneme left and one right and we extract the corresponding utterance with the help of forced alignment. We go from no context to context length 7 (both right and left) and we compare to the results obtained with the original utterances (full context).
4) How do our results compare to the state-of-the-art in MDD? In order to compare our results to state-of-the-art, we tested our best method (GOP-X-AF-SD) on the Speechocean762 dataset, and we also include FGOP, that is, alignment-free GOP features. Moreover, we compare our new implementations that solve the numerical issues (denoted as -numerical) and the normalized versions of both the scalar GOP and the GOP features. Because the focus is on comparing between different GOP and GOP-features definitions, we keep the models for MDD simple in our state-of-the-art comparison. In particular, we limit our tests to polynomial regression, support vector regression (SVR) and GOPT as in [GOPT]. The PCC is computed between the model’s output value and the the label. For the sake of fair comparison, we preserve all the details for evaluations as the baseline papers, e.g, whether to round the hypothesis before calculating the PCC; using polynomial order two for polynomial regression; applying the Radial Basis Function (RBF) kernel for the SVR models etc. Due to relatively high variance of the GOPT model, same as in [GOPT], we run the training of GOPT five times with random initialization, selecting the best model for each run and then averaging the results. Following the same idea in [cao23_interspeech] to reduce the variance of the GOPT model during training, we limit the gradients to flow from phoneme-only loss instead of from all multi-aspects losses.
@@ V-A Method comparison
Table I shows the results comparing the different methods on the CMU Kids data both for simulated and real errors. AUC values are reported together with 95% confidence intervals computed according to [HanleyAndMcNeil1982ROC-CI]. The methods based on models trained with CTC outperform both methods based on models trained with CE (GOP-DNN-Avg and GOP-CE-Avg). GOP-CTC-AF-S, that is, the alignment-free method restricted to substitution errors is significantly better on the simulated errors whereas GOP-CTC-AF-SD, that also allows deletion errors is significantly better on the real errors. This can be explained by the fact that simulated errors are produced exclusively by substitution, but real errors may include deletions as well. Including insertions (GOP-CTC-AF-SDI) does not seem to help for this dataset.
@@ V-B Performance versus peakiness
Table II shows the results of testing the effect of peakiness of the acoustic models on the real errors of the CMU-kids. All the models were trained up to the 8000th checkpoint. The left part of the table reports blank coverage (BC) as a measure of POT and conditional entropy (ConEn) as a measure of POS, as well as phone error rates. The right part of the table reports MDD results. The models were ordered in the table in increasing order of peakiness over time (POT).
@@
The model trained with EsCTC has the highest PER, a similar POT compared to CTC but a much lower POS (higher ConEn). This is in line with the experiments in [CTC-ME], because the model makes the strong assumption that the labels should be uniformly distributed over time which is not always true for speech.
@@
As we expected, the model trained with CE has the lowest BC because it is trained frame-wise according to baseline alignment. Surprisingly, the ConEn of the CE-trained model is still lower than EnCTC implying that the distribution of possible alignment paths are more concentrated for CE-trained model.
@@
The experimental results for MDD are shown in the right side of Table II in terms of AUC and 95% confidence intervals. When using the standard GOP definition (GOP-X-Avg), the CE model performs the best. This is expected because GOP-X-Avg uses the same baseline alignment that was used to train the CE model. The CTC and EnCTC models are comparable. We believe that this is because, although the models have different degrees of peakiness, their activation is still within the segments defined by the baseline alignment. On the contrary, the AUC of GOP-EsCTC-Avg is particularly low, most probably because the model activations do not necessarily follow standard speech segmentation. It is therefore possible that GOP is evaluated, in this case, based on activations other than the target phoneme. This interpretation is also supported by the fact that the use of the self-aligned definition (GOP-EsCTC-SA) gives results that are comparable with the model trained with CE. However, GOP-X-Avg is clearly outperformed by the proposed GOP-X-SA and GOP-X-AF methods that will be analyzed next.
@@
In general, the proposed GOP-X-SA is always superior to GOP-X-Avg, suggesting that this method is able to focus on the regions corresponding to the relevant model activations. It is interesting to note that this phenomenon can be observed even for the models trained with CE, for which the mismatch between model activations and baseline alignment should be minimal. Among the different models, CTC training performs the best when using GOP-X-SA. Our interpretation is that the self-aligned definition of GOP (GOP-X-SA) prefers to work with peaky models (both POT and POS) such as the model trained with CTC. Comparing to the standard definition (GOP-X-Avg), this method makes it possible to take advantage of the highly-discriminative model which is often good for recognition. This interpretation is also supported by the slight degradation in AUC results compared to standard CTC for the EnCTC models which are less peaky and less performant in terms of PER. Exceptionally, the behavior of the model trained with EsCTC has by far the worst phone recognition performance. However, the model’s peakiness (with the second highest BC and second lowest ConEn) could still make it comparable with CE in MDD performance when combined with GOP-X-SA.
@@
Finally, the last column of the table reports the results with the proposed alignment-free method (GOP-X-AF). This method obtains the best overall performance for MDD by a significant margin. Furthermore, according to the confidence intervals it is not possible to determine which acoustic model performs the best with this method because the results obtained with CTC and EnCTC are not statistically different. This is in accordance with our hypothesis that GOP-X-AF would be less affected by the peakyness over time of the acoustic models, compared to methods that rely on alignment. We can also observe that, by considering all possible alignments, GOP-X-AF is able to extract much more useful information for MDD not only with respect to the external alignment (GOP-X-Avg) but also with respect to an alignment that is matched to the activations of the acoustic model (GOP-X-SA).
@@
Figure 4: AUC versus context length for simulated and real errors on the CMU Kids data. The shaded area shows 95% confidence intervals computed with [HanleyAndMcNeil1982ROC-CI].
V-C GOP-AF and context length
@@
By definition, the alignment-free GOP method that we have proposed (GOP-AF) is computed considering contributions from the entire utterance, even if those contributions are weighted by how likely it is that each part of the utterance belongs to the target phone. A reasonable question to ask is how the length of left context ($L_{\text{L}}$) and right context ($L_{\text{R}}$) affect the pronunciation assessment results.
@@
Figure 4 displays the AUC results on CMU-Kids where we have varied the number of phonemes in $L_{\text{L}}$ and $L_{\text{R}}$ from 0 to 7 resulting in a total context length from 0 to 14 in increments of 2. We also report results with the original utterances (full context). In this case, the length of the context depends on the specific utterance, but it is always greater than 14. The left plots simulated errors, whereas the right plots real errors.
@@
For simulated errors, the AUC values for different context lengths remain at a high level, and the impact of context length can be neglected. The results show that GOP-CTC-AF is robust to different context length in the ideal case.
@@
Also for real errors, the dependency on context length is usually under the variability described by the 95% confidence intervals. The only exception is when we reduce the context to zero, which corresponds to the self-aligned GOP definition (GOP-SA). This confirms the superiority of the alignment-free method that is able to take advantage of the context to assess the pronunciation of the target phoneme. Using context lengths between 8 and 14 is visually indistinguishable from using the full context length. This suggests that the information needed to assess the pronunciation of the target phoneme is relatively local in the utterance.
TABLE III: Speechocean762, reported with average F1 and Pearson correlation coefficients
@@
GOP-TDNN [cao24b_interspeech] poly. reg. 0.361$\pm$0.008
GOP-CE-AF-SD poly. reg. N/A
GOP-CE-AF-SD-numerical poly. reg. 0.314$\pm$0.008
GOP-CE-AF-SD-Norm poly. reg. 0.332$\pm$0.008
GOP-CTC-AF-SD [cao24b_interspeech] poly. reg. 0.433$\pm$0.007
GOP-CTC-AF-SD-numerical poly. reg. 0.450$\pm$0.006
GOP-CTC-AF-SD-Norm poly. reg. 0.449$\pm$0.006
FGOP-TDNN [speechocean762] SVR 0.441$\pm$0.007
FGOP-CTC-AF [cao24b_interspeech] SVR 0.568$\pm$0.006
FGOP-CTC-AF-numerical SVR 0.580$\pm$0.006
FGOP-CTC-AF-Norm SVR 0.581$\pm$0.006
FGOP-TDNN [GOPT] GOPT 0.605$\pm$0.002
FGOP-CTC-AF [cao24b_interspeech] GOPT 0.618$\pm$0.002
FGOP-CTC-AF-numerical GOPT 0.646$\pm$0.002
FGOP-CTC-AF-Norm GOPT 0.648$\pm$0.002
@@ V-D Comparison to state-of-the-art
In order to compare the performance of our best method (GOP-X-AF-SD) with respect to the state-of-the-art, we report results on the SpeechOcean762 dataset in Table III. The table compares different scalar definitions of GOP with a simple polynomial regression MDD model. It also includes results with GOP feature vectors in combination with SVR or GOPT MDD models.
Similar to the previous experiments on CMU-kids, for CE-trained model, we find that the methods are less performant when comparing to the CTC-trained model. Results for the method “GOP-CE-AF-SD” are missing due to numerical problems. These were solved in the new implementation (“GOP-CE-AF-SD-numerical”). The results are further improved using the definition that normalizes GOP by the estimated length of the target segment (“GOP-CE-AF-SD-Norm”).
In all cases, when combined with a CTC-trained model, our alignment-free methods (GOP-CTC-AF, and FGOP-CTC-AF) outperform the baseline methods that use traditional GOP with the Time-Delayed-Neural-Network (TDNN) acoustic models by a significant margin. We also see improvements when using the implementation which solves the numerical issues compared to our previous results. As we expected, the extra normalization steps for the methods are less significant for CTC-trained acoustic models than CE-trained ones. The best combination of our method is “FGOP-CTC-AF-Norm” which has $5.01\%$ relatively higher PCC than previously reported. To the best of our knowledge, the highest phoneme-level PCC obtained for this dataset is $0.693$ [chao23_interspeech]. However, their work uses a more complex model than GOPT, more heavy feature engineering and more sophisticated loss functions that take all the linguistic aspects into account.

### v3

@@ IV-C Pre-training and fine-tuning of the acoustic models
The acoustic models ``DNN'' (CMU Kids baseline) and ``TDNN'' (speechocean762 baseline) are trained from scratch using the Kaldi recipes666./egs/librispeech/s5/local/{nnet2/run_5a_clean_100, nnet3/run_tdnn}.sh. All the other models are based on the wav2vec2.0-large-xlsr-53 [6] which is available on Huggingface [^7] https://huggingface.co/blog/fine-tune-xlsr-wav2vec2 and have a transformer-based structure [1]. This model, with the feature_extractor frozen, was fine-tuned on the Librispeech ``train-100-clean'' set according to the corresponding loss functions:
@@
  • EnCTC: CTC with entropy regularization [26]
  • EsCTC: CTC with equal space constraint [26]
Fine-tuning was performed for at most 10 epochs or by early stopping based on the validation loss, with learning rate 1e-4 and batch size 32. The same model was used for the CMU Kids and speechocean762 experiments.
@@ IV-D Evaluation
The evaluation on CMU Kids is based on the canonical transcription for each utterance. Each phone in the canonical transcription is marked as mispronounced if it deviates from the phonetic transcription from the data set. Then we compute the receiver operative characteristic curve (AUC-ROC) according to various GOP scores, where correctly detecting a mispronunciation is a ``true positive''. We opt for this threshold-free criterion because we want a robust method to evaluate the performance of our GOP definitions that is not dependent on a specific threshold chosen to perform the MDD task. This also allows us to use all the data for testing because we do not need to set aside a training set for threshold optimization. Moreover, AUC-ROC is considered to be robust to label bias which is the case for our data. We compute AUC-ROC for each phoneme category and report the arithmetic mean.
For the MDD-oriented dataset speechocean762 the task is to predict three different classes. In this case, we train and evaluate several typical MDD models using the standard training and test splits. We follow the recommendations of using Pearson Correlation Coefficient (PCC) from the dataset's paper [40] as MDD performance metric.
To give a reference on the quality of the different acoustic models in this study, we also report on phoneme recognition results measured with phoneme error rate (PER). Beam-search decoding without language models is used for recognition.
@@ IV-E Experiments
The set of experiments we include in this paper aims at answering the following questions:
1) Which GOP definition results in the best MDD performance? In this case we test GOP-X-Avg, GOP-X-SA, and the different variants of GOP-X-SF (GOP-X-SF-SDI, GOP-X-SF-SD, GOP-X-SF-S) where ``X'' corresponds to acoustic models trained with different loss functions and architectures as explained in Sec.IV-C.
2) How are the results dependent on the peakiness of the ASR models used for the assessment? In this case we test the different GOP definitions in combination with models that are trained with loss functions that are specifically defined to control peakiness: CTC, EnCTC, EsCTC and CE. We use our best method GOP-SF-SD as segmentation-free version in this experiment and due to numerical problem with CE-trained models, we use the ``numerical'' implementation as discussed in Sec.III-B.
3) How do the results for GOP-SF depend on the length of the context around the target phoneme? In this case we create utterances by cropping the original recordings and gradually increasing the context around the target phoneme. Each time we add one phoneme left and one right and we extract the corresponding utterance with the help of forced alignment. We go from no context to context length 7 (both right and left) and we compare to the results obtained with the original utterances (full context).
4) How do our results compare to the state-of-the-art in MDD? In order to compare our results to the state-of-the-art, we tested our best method (GOP-X-SF-SD) on the speechocean762 dataset, and we also include FGOP, that is, segmentation-free GOP features. Moreover, we compare our new implementations that solve the numerical issues (denoted as -numerical) and the normalized versions of both the scalar GOP and the GOP features. Because the focus is on comparing between different GOP and GOP-features definitions, we keep the models for MDD simple in our state-of-the-art comparison. In particular, we limit our tests to polynomial regression, support vector regression (SVR) and GOPT as in [13]. The PCC is computed between the model's output value and the true label. For the sake of fair comparison, we preserve all the details for evaluations as the baseline papers, e.g, whether to round the output before calculating the PCC; using polynomial order two for polynomial regression; applying the Radial Basis Function (RBF) kernel for the SVR models etc. Due to relatively high variance of the GOPT model, same as in [13], we run the training of GOPT five times with random initialization, selecting the best model for each run and then averaging the results. Following the same idea in [3] to reduce the variance of the GOPT model during training, we limit the gradients to flow from phoneme-only loss instead of from all multi-aspects losses.
@@ V-A Method comparison
Table I shows the results comparing the different methods on the CMU Kids data both for simulated and real errors. AUC values are reported together with 95% confidence intervals computed according to [15]. The methods based on models trained with CTC outperform both methods based on models trained with CE (GOP-DNN-Avg and GOP-CE-Avg). GOP-CTC-SF-S, that is, the segmentation-free method restricted to substitution errors is significantly better on the simulated errors whereas GOP-CTC-SF-SD, that also allows deletion errors is significantly better on the real errors. This can be explained by the fact that simulated errors are produced exclusively by substitution, but real errors may include deletions as well. The standard evaluation criterion used here only considers phones in the canonical transcription. This means that insertion errors only have an effect on the evaluation if they modify the pronunciation of the neighbouring canonical phones. This does not allow to show the full potential of GOP-CTC-SF-SDI, that in our tests performs worse than GOP-CTC-SF-SD. However, defining an ad-hoc evaluation criterion is outside the scope of this study.
@@ V-B Performance versus peakiness
Table II shows the results of testing the effect of peakiness of the acoustic models on the real errors of the CMU Kids. All the models were trained up to checkpoint 8000. The left part of the table reports blank coverage (BC) as a measure of POT and conditional entropy (ConEn) as a measure of POS, as well as phone error rates. The right part of the table reports MDD results. The models were ordered in the table in increasing order of peakiness over time (POT).
@@
The model trained with EsCTC has the highest PER, a similar POT compared to CTC but a much lower POS (higher ConEn). This is in line with the experiments in [26], because the model makes the strong assumption that the labels should be uniformly distributed over time which is not always true for speech.
@@
As we expected, the model trained with CE has the lowest BC because it is trained frame-wise according to baseline segmentation. Surprisingly, the ConEn of the CE-trained model is still lower than EnCTC implying that the distribution of possible alignment paths are more concentrated for CE-trained model.
@@
The experimental results for MDD are shown in the right side of Table II in terms of AUC and 95% confidence intervals. When using the standard GOP definition (GOP-X-Avg), the CE model performs the best. This is expected because GOP-X-Avg uses the same baseline segmentation that was used to train the CE model. The CTC and EnCTC models are comparable. We believe that this is because, although the models have different degrees of peakiness, their activation is still within the segments defined by the baseline segmentation. On the contrary, the AUC of GOP-EsCTC-Avg is particularly low, most probably because the model activations do not necessarily follow standard speech segmentation. It is therefore possible that GOP is evaluated, in this case, based on activations other than the target phoneme. This interpretation is also supported by the fact that the use of the self-aligned definition (GOP-EsCTC-SA) gives results that are comparable with the model trained with CE. However, GOP-X-Avg is clearly outperformed by the proposed GOP-X-SA and GOP-X-SF methods that will be analyzed next.
@@
In general, the proposed GOP-X-SA is always superior to GOP-X-Avg, suggesting that this method is able to focus on the regions corresponding to the relevant model activations. It is interesting to note that this phenomenon can be observed even for the models trained with CE, for which the mismatch between model activations and baseline segmentation should be minimal. Among the different models, CTC training performs the best when using GOP-X-SA. Our interpretation is that the self-aligned definition of GOP (GOP-X-SA) prefers to work with peaky models (both POT and POS) such as the model trained with CTC. Comparing to the standard definition (GOP-X-Avg), this method makes it possible to take advantage of the highly-discriminative model which is often good for recognition. This interpretation is also supported by the slight degradation in AUC results compared to standard CTC for the EnCTC models which are less peaky and less performant in terms of PER. Exceptionally, the behavior of the model trained with EsCTC has by far the worst phone recognition performance. However, the model's peakiness (with the second highest BC and second lowest ConEn) could still make it comparable with CE in MDD performance when combined with GOP-X-SA.
@@
Finally, the last column of the table reports the results with the proposed segmentation-free method (GOP-X-SF). This method obtains the best overall performance for MDD by a significant margin. Furthermore, according to the confidence intervals it is not possible to determine which acoustic model performs the best with this method because the results obtained with CTC and EnCTC are not statistically different. This is in accordance with our hypothesis that GOP-X-SF would be less affected by POS of the acoustic models, compared to methods that rely on segmentation. We can also observe that, by considering all possible segmentations, GOP-X-SF is able to extract much more useful information for MDD not only compared to the external segmentation (GOP-X-Avg) but also compared to an alignment that is matched to the activations of the acoustic model (GOP-X-SA).
@@
Figure 4: AUC versus context length for simulated and real errors on the CMU Kids data. The shaded area shows 95% confidence intervals computed with [15].
V-C GOP-SF and context length
@@
By definition, the segmentation-free GOP method that we have proposed (GOP-SF) is computed considering contributions from the entire utterance, even if those contributions are weighted by how likely it is that each part of the utterance belongs to the target phone. A reasonable question to ask is how the length of left context ($L_{\text{L}}$) and right context ($L_{\text{R}}$) affect the pronunciation assessment results.
@@
Figure 4 displays the AUC results on CMU Kids where we have varied the number of phones in $L_{\text{L}}$ and $L_{\text{R}}$ from 0 to 7 resulting in a total context length from 0 to 14 in increments of 2. We also report results with the original utterances (full context). In this case, the length of the context depends on the specific utterance, but it is always greater than 14. The left plots simulated errors, whereas the right plots real errors.
@@
For simulated errors, the AUC values for different context lengths remain at a high level, and the impact of context length can be neglected. The results show that GOP-CTC-SF is robust to different context length in the ideal case.
@@
Also for real errors, the dependency on context length is usually under the variability described by the 95% confidence intervals. The only exception is when we reduce the context to zero, which corresponds to the self-aligned GOP definition (GOP-SA). This confirms the superiority of the segmentation-free method that is able to take advantage of the context to assess the pronunciation of the target phoneme. Using context lengths between 8 and 14 is visually indistinguishable from using the full context length. This suggests that the information needed to assess the pronunciation of the target phoneme is relatively local in the utterance.
TABLE III: speechocean762, reported with Pearson correlation coefficients
@@
[…]
GOP-CE-SF-SD poly. reg. N/A
[…]
@@ V-D Comparison to state-of-the-art
In order to compare the performance of our best method (GOP-X-SF-SD) with respect to the state-of-the-art, we report results on the speechocean762 dataset in Table III. The table compares different scalar definitions of GOP with a simple polynomial regression MDD model. It also includes results with GOP feature vectors in combination with SVR or GOPT MDD models.
Similarly to the previous experiments on CMU Kids, we find that the methods relying on models trained with the CE loss are less performant than those trained with the CTC loss. Results for the method ``GOP-CE-SF-SD'' are missing due to numerical problems. These were solved in the new implementation (``GOP-CE-SF-SD-numerical''). The results are further improved using the definition that normalizes GOP by the estimated length of the target segment (``GOP-CE-SF-SD-Norm'').
In all cases, when combined with a CTC-trained model, our segmentation-free methods (GOP-CTC-SF, and FGOP-CTC-SF) outperform the baseline methods that use traditional GOP with the Time-Delayed-Neural-Network (TDNN) acoustic models by a significant margin. We also see improvements when using the implementation which solves the numerical issues compared to our previous results. As we expected, the extra normalization steps for the methods are less significant for CTC-trained acoustic models than CE-trained ones. The best combination of our method is ``FGOP-CTC-SF-Norm'' which has $5.01\%$ relatively higher PCC than previously reported. We also tested this method on the split of the speechocean762 test set that separates children (age $\leq 11$) from adults (age >11>11). The corresponding PCCs are $0.635\pm 0.005$ (children) and $0.651\pm 0.004$ (adults). The slight gap between child and adult result may be reduced by adapting the acoustic models to child speech.
At the time of writing and to the best of our knowledge, the highest phoneme-level PCC obtained for the speechocean762 dataset is $0.693$ [5]. The relatively limited PCC improvements of this model with respect to ours are, however, obtained at the cost of higher computational complexity, heavier feature engineering and more sophisticated loss functions that take all the linguistic aspects into account. For this reason, we believe that our method is still preferable for practical real-world applications.
Finally, we could not find results on pure end-to-end models for MDD on the speechocean762 data. This is probably due to the limited amount of training data which prevents the proper training of those models.